A power of 4.5 kW must be transmitted from a pulley of an effective diameter of 0.25 m. The pulley is running at 1200 RPM. The angle of contact is 160° and the groove angle is 40°. The coefficient of friction is 0.2. If the allowable tension in each belt is 100 N, calculate the number of V-belts that will be required to transmit the power.


1
Expert's answer
2021-10-10T09:33:53-0400

Solution;

Given;

Power transmitted,P=4.5kW

Effective diameter,"d_{eff}=0.25m"

Speed of pulley,N="1200" r.p.m

Angle of contact,"\theta" =160°

Maximum tension,"T_1" =100N

Co-efficient of friction,"\mu" =0.2

Groove angle,2"\beta" =40°; "\beta" =20°

Contact angle,"\theta" ="\frac{π×160}{180}=2.792rad"

By using the relation;

"2.3log(\frac{T_1}{T_2})=\mu×\theta×cosec\beta"

"log(\frac{T_1}{T_2})=\frac{0.2×2.792×cosec(20°)}{2.3}"

"log(\frac{T_1}{T_2})=0.7098"

"\frac{T_1}{T_2}=e^{0.709}" "=2.0336"

"T_2=\frac{T_1}{2.0336}=\frac{100}{2.0336}"

"T_2=49.172N"

Torque transmitted;

"T=(T_1-T_2)×\frac{d_{eff}}{2}"

"T=(100-49.172)\u00d70.125"

"T=6.3535Nm"

Power transmitted by a single belt;

"P_s=\frac{2πNT}{60}"

"P_s=\frac{2π×1200×6.35}{60}" ="798.40W"

Number of belts;

"n=\frac{P}{P_s}=\frac{4500}{798.4}=5.6"

"n=5.636\approx6"

Use 6 belts.

Answer;

6 belts.